Problem 14-9 Zero-coupon bonds [LO14-2] On January 1, 2018, Darnell Window and Pane issued $18.5 million of 10-year, zero-coupon bonds for $6,515,413. (FV of $1. PV of $1 FVA of $1, PVA of $1, FVAD of $1 and PVAD of $1 (Use appropriate factor(s) from the tables provided.) Required 2. Determine the effective rate of interest. 1. & 3.to 5. Prepare the necessary journal entries. Complete this question by entering your answers in the tabs General Interest rateJournal Determine the effective rate of interest. nterest rate K Interest rate General Journal>

Complete this question by entering your answers in the tabs below General Journal Interest rate Prepare the necessary journal entries. (If no entry is required for a transaction/event, select No journal entry required in the first account field. Round your intermedaite calculation and final answers to nearest whole dollar.) View transaction list 1 Record the issuance of the bonds. 2 Record annual interest expense at December 31, 2018. / 3 Record annual interest expense at December 31, 2019. 4 Record the payment at the bonds maturity. Credit
Answer #1

2)

Price of Bond = $6515413

Maturity (n) = 10 years

Face value of Bond = $18500000

Let effective rate of interest be r.

Price of Bond = Face value of Bond/ (1+r) ^n

$6515413 = $18500000/ (1+r) ^10

(1+r)^10 = $18500000/$6515413

(1+r)^10 = 2.8394

1+r = (2.8394) ^1/10

1+r = 1.11

r = 1.11 – 1 = 0.11 i.e. 11%

Effective rate of interest is 11%.

1) & (3 to 5)

Sl. NO.

Account Titles

Debit($)

Credit($)

1)

Cash

6515413

   Bond Payable

6515413

(Issuance of Bond)

2)

Interest Expense

716695

(6515413*11%)

    Bond Payable

716695

(Annual interest expense as at December 31,2018)

3)

Interest Expense

795532

[(6515413+716695)*11%]

    Bond Payable

795532

(Annual interest expense as at December 31,2019)

4)

Bond Payable

18500000

   Cash

18500000

(Payment at Bond maturity)